# Liouville Flow Importance Sampler

### v1

@@ 1.1 Related Work
Numerous Monte Carlo (MC) techniques have been devised to address this challenge over the past 70 years. Some notable examples of these MC methods include the landmark Metropolis Markov chain MC (MCMC, (Metropolis et al., 1953)), Hybrid or Hamiltonian MC (HMC, Duane et al. (1987); Neal (2012)), and the Zig-Zag Sampler (Bierkens et al., 2019). These approaches are often limited to low-dimensional problems due to slow convergence rate, and problems with a single mode as the samplers are prone to be trapped at a local mode. Annealed Importance Sampler (AIS, Neal (2001)) and Sequential Monte Carlo (SMC, (Del Moral et al., 2006)) were developed for sampling challenging multi-modal distributions and are considered the state-of-the-art MC methods for such problems. Variational Inference (VI, Wainwright & Jordan (2008)) is an alternative approach, which turns the sampling problem into an optimization one by fitting an easy-to-sample parametric distribution to $\nu$, often by minimizing the reverse (or exclusive) Kullback–Leibler (KL) divergence. The choice of the reverse KL is commonly thought to be due to computational convenience but can be validated by an information theoretical argument of optimal information processing of Bayes’ theorem (Zellner, 1988). For VI, it is popular to adopt mean-field approximation (Wainwright & Jordan, 2008) or a Normalizing Flow (VI-NF, Rezende & Mohamed (2015)) as the modeling parametric distribution. More recently, a plethora of neural network (NN) based algorithms, for example, neural network gradient HMC (Li et al., 2019), Annealed Flow Transport Monte Carlo (AFTMC, Arbel et al. (2021a)), Path-Integral Sampler (PIS, Zhang & Chen (2022)), adaptive MC with NF Gabrié et al. (2022), Continual Repeated AFTMC (CR-AFTMC, Matthews et al. (2022)), and Denoising Diffusion Sampler (DDS, (Vargas et al., 2023a)). For estimating the marginalized likelihood, simple MCMC methods require augmented algorithms (e.g., Green (1995)) or additional processing of the derived samples (e.g., Robert & Wraith (2009); Pajor (2017)). SIS and SMC can estimate the marginalized likelihood and are considered the gold-standard MC estimates. Among the NN-based algorithms, AFTMC, CR-AFTMC, PIS, and DDS are capable of estimating the marginalized likelihood. Llorente et al. (2023) provides a more comprehensive review of methods estimating the marginalized likelihood.
@@ 2.2 Learning velocity field for sampling
We remark that our methodology significantly differs from the Flow Matching approach (Lipman et al., 2023; Tong et al., 2024). Flow Matching trains a continuous-time normalizing flow by matching it to a known velocity field. This technique is specifically designed for training Generative Models using samples of a data distribution. In contrast, our task involves learning the normalizing flow using a prescribed path of annealed distributions, $\tilde{\rho}_{\ast}(x,t)$. In this scenario, we do not have access to the samples of the target distribution nor a target velocity field for matching: the problem would have been solved if we had the samples of the target distribution, and we could have simply driven the samples by a known target velocity field. Instead, our knowledge is limited to the fact that the velocity field at each of the samples must satisfy Eq. (5). Our goal is to solve for the unknown $v(x,t)$ in Eq. (5), rather than attempting to match a target flow.
@@ 3 Numerical Experiments
After training, we performed 30 independent samplings (each with $S=2000$ samples, different from those used in training) using the learned velocity fields. Following the recent series of papers (Arbel et al., 2021a; Zhang & Chen, 2022; Vargas et al., 2023a), we focus on assessing sample qualities by the estimation of log normalization constant $\log\mathcal{Z}$. For low dimensional ($D\leq 10$) problems, we also compare the sliced Wasserstein distance $W_{p}$ (Bonneel et al., ) to ground-truth samples with $p=2$ as an additional metric. For all the methods with importance sampling, we used the weighted samples for computing these statistical quantities.
@@ 3.1 Testing problems
Mode-separated Gaussian mixture (type-1): An illustrative model of $D=2$ Gaussian mixture distribution with separated modes. We consider a similar challenging distribution as in PIS (Zhang & Chen, 2022): a mixture of nine Gaussian distributions centered at the grid $\{-1,0,1\}^{2}$, and each Gaussian has variance 0.012.
@@ 3.2 Results
Tables 1 and 2 show the performance of different methods on type-1 and type-2 problems respectively. We visualize the samples and weight distributions of type-1 problems in Fig. 2. Comprehensive results of different experimental settings can be found in Appendix D. Our numerical results showed that LFIS is capable of generating competitive samples to existing methods. For two type-1 applications, we observed reasonable estimates of $\log\mathcal{Z}$. Ground-truth (GT) samples can be generated for these two test problems, and the samples generated by LFIS have the lowest sample-to-GT-sample Wasserstein-2 distance. We remark that LFIS performs even better than the conventionally regarded gold-standard SMC with the same number of scales ($T=256$). For Effective Sample Size (ESS) (Liu, 2008), LFIS also shows competitive performance among all methods studied. Note that both SMC and AFT perform resampling when ESS is lower than a certain threshold (0.98 for SMC and 0.3 for AFT (Arbel et al., 2021a)), while LFIS, DDS, PIS do not resample. Even without resampling, LFIS can still achieve high ESS comparable to SMC with a high resampling threshold. The samples generated by LFIS also provide more uniform coverage over the different modes of the Gaussian mixture and the tails of funnel distribution compared to all other methods. In addition, the weight distributions of LFIS in these two problems are significantly narrower than other methods, explaining the higher ESS and more accurate estimation. For type-2 problems, LFIS is the only method delivering estimates of $\log\mathcal{Z}$ that are consistent with the gold-standard SMC (with $T=1024$) on all problems.
@@ 4 Discussion
VI-NF (Rezende & Mohamed, 2015) uses a normalizing flow as the modeling distribution, which is parametrized by variational inference. LFIS shares many similar features of VI-NF when a Neural ODE (Chen et al., 2018) is used as the flow. In this case, both our proposition and VI-NF draw samples from simple initial distributions ($\mu$ in our case; the base distribution in VI-NF) and use the NN-modeled flow to transport the samples to a future time. Both methods leverage the computable density function along the sample trajectory, a unique feature of flow-based models (see Eq. (8)). The key difference between our proposition and VI-NF is the learning target. VI-NF is an end-to-end approach, that the loss function only depends on matching the final distribution at the end of the flow to the target distribution. In contrast, LFIS aims to capture $\tilde{\rho}_{\ast}(t)$, $t\in\left[0,1\right]$. We hypothesized that providing the whole evolution $\tilde{\rho}_{\ast}(t)$, instead of only the final $\tilde{\nu}=\rho_{\ast}(1)$, could improve the learning [^3] As the same philosophy applies to AIS and SMC (analogous to LFIS) to simple MCMC-type inference (analogous to VI-NF).. Moreover, VI-NF aims to directly optimize the reverse Kullback–Leibler (KL) divergence. In contrast, our approach first establishes the PDE (5) for the optimal flow which perfectly matches the evolution of the target distribution. Then, we parametrize the neural flow model by minimizing its discrepancy to the evolutionary equation, an approach similar to the Physics Informed Neural Networks (Raissi et al., 2019; Karniadakis et al., 2021). Because VI-NF has been shown to perform worse than AFTMC (Arbel et al., 2021a), we did not include VI-NF in our quantitative comparison in Sec. 3.
@@
We conclude the manuscript by listing the potential limitations of LFIS. (1) LFIS cannot handle non-differentiable density function $\tilde{\nu}(x)$. For example, bounded uniform priors. Theoretically, these singularities can be handled by inserting sources at the discontinuities, a direction that merits future developments. A practical solution to address the challenge is to consider the non-differentiable density function as the limit of differentiable $\tilde{\rho}_{\ast}(x,t)$. For example, using sigmoid functions $\tilde{\nu}(x,t)=\sigma(tx/(1-t))$ which converge to step functions $\tilde{\nu}(x)={\Theta}(x)$. (2) When the flow satisfying Eq. (5) is too complex, LFIS’ approach may require a more expressive NN than end-to-end DDS and PIS. This can result in a more resource-intensive training process, but we speculate that it might offer a tradeoff in terms of improved accuracy. (3) LFIS is more memory-demanding as it requires $\nabla_{x}S_{\ast}$ in the equation, and cross terms like $\partial_{\theta}\partial_{x}v_{\theta}(x,t)$ in the optimization step. These terms are currently evaluated by memory-demanding autograd. (4) As an integrator and with the currently adopted explicit scheme, LFIS does not perform well when $T$ is small due to the error induced by a finite time-step (see Fig. 7). This may result in a higher training cost. Optimizing LFIS using higher-order integration schemes and an interpolation of the neural flow along the time domain merits future research.
Code availability. The Feynmann Center for Innovation of Los Alamos National Laboratory is currently reviewing our codes for unlimited release. Once approved, we will deposit the codes at https://github.com/lanl/LFIS.
Acknowledgments. The authors acknowledge continual support from Laboratory Directed Research and Development (LDRD). YT was supported by LDRD project “Accelerated Dynamics Across Computational and Physical Scales” (220063DR), NP was supported by LDRD project “Learning Uncertainties In Coupled-Physics Models via Operator Theory” (20230254ER), and YTL was supported by LDRD project “Diffusion Modeling with Physical Constraints for Scientific Data” (20240074ER). YTL sincerely thanks Prof. A. Doucet for several insightful email exchanges that partially inspired this work.
Broader impact. This manuscript presents work whose goal is to advance the field of statistical sampling and machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix B Importance Sampling for Imperfect Neural Networks
We put a stringent condition that $\rho_{\theta}$ has to dominate $\rho_{\ast}$ for any arbitrary function $F$. The condition is often relaxed to “$\rho_{\theta}(x)>0$ where $F(x)\rho_{\ast}(x)>0$ in the importance sampling literature (e.g., Tokdar & Kass (2010)), which requires knowledge of the function $F$.
@@ D.1 Additional details of the test distributions
For the sampling problem in the latent space of VAE, we took to decoder $p_{\theta}(x|z)$ of the pre-trained VAE in Arbel et al. (2021a) and used it to construct the likelihood function with binary cross-entropy loss. In Fig. 3, we demonstrate the test image and the reconstructed image using the pre-trained decoder $p_{\theta}(x|z)$ from samples generated by LFIS.
@@ D.3 Neural network architecture
It is difficult to make a direct comparison of NN structures and numbers of parameters between the discrete-time and continuous-time models. For LFIS, the total number of parameters will increase with the total number of steps. For DDS and PIS, the number of parameters will increase with the number of channels used for time-encoding increasing. Overall, when LFIS uses the same number of time steps as the time-encoding in DDS and PIS, the total number of NN parameters are very similar, thus causing no additional memory footprint.
@@ D.4 Additional details of the training procedure
The general training procedure is summarized in Algorithm 1. For each discretized time step $k$, at each training epoch, a new minibatch (B) of samples is generated by calling the GenerateSamples function. This step is crucial to provide good coverage of the sampling space. However, for low-dimensional problems, repeatedly calling the GenerateSamples function for each minibatch could slow down the training process, because the initial samples need to be passed through all the previously trained flow $v_{\theta_{\ast}^{(l)}}$. To accelerate the training of low-dimensional sampling problems ($D<10$), we skip the generation of new samples at each epoch. Instead, we use a subset of the $N$ samples used for estimating $\langle\partial_{t}\log\tilde{\rho_{\ast}}\rangle_{\ast}$ for training. When $N$ is large enough to provide good coverage of the sampling space, this simplified training procedure could provide a noticeable acceleration of learning the Liouville flow (add numbers).
@@ D.9 Choice of total time steps and tempered scales
32 $-0.064\pm 0.065$ $-0.43\pm 0.050$ $-220.70\pm 2.23$ $-112.28\pm 0.68$ $-133.36\pm 2.08$ $470.13\pm 5.57$
$\widehat{\log\mathcal{Z}}$ 64 $-0.019\pm 0.060$ $-0.27\pm 0.066$ $-110.58\pm 0.39$ $-112.04\pm 0.38$ $-108.94\pm 0.30$ $489.31\pm 3.20$
128 $-0.007\pm 0.066$ $-0.23\pm 0.057$ $-110.07\pm 0.31$ $-111.67\pm 0.26$ $-108.60\pm 0.31$ $505.26\pm 5.56$
256 $0.004\pm 0.059$ $-0.11\pm 0.075$ $-109.95\pm 0.25$ $-111.63\pm 0.33$ $-108.48\pm 0.30$ $505.43\pm 3.00$
32 $0.006\pm 0.005$ $-0.29\pm 0.023$ $-215.64\pm 0.87$ $-111.42\pm 0.01$ $-130.45\pm 0.233$ $463.40\pm 2.61$
$\log\hat{\mathcal{Z}}$ 64 $0.002\pm 0.003$ $-0.16\pm 0.028$ $-109.98\pm 0.01$ $-111.58\pm 0.006$ $-108.36\pm 0.011$ $484.54\pm 2.01$
128 $-0.0001\pm 0.004$ $-0.20\pm 0.011$ $-109.95\pm 0.01$ $-111.59\pm 0.004$ $-108.37\pm 0.008$ $504.32\pm 1.56$
256 $-0.0002\pm 0.004$ $-0.07\pm 0.003$ $-109.99\pm 0.08$ $-111.60\pm 0.006$ $-108.38\pm 0.009$ $505.53\pm 0.95$

### v2

@@ 1.1 Related Work
Numerous Monte Carlo (MC) techniques have been devised to address this challenge over the past 70 years. Some notable examples of these MC methods include the landmark Metropolis Markov chain MC (MCMC, (Metropolis et al., 1953)), Hybrid or Hamiltonian MC (HMC, Duane et al. (1987); Neal (2012)), and the Zig-Zag Sampler (Bierkens et al., 2019). These approaches are often limited to low-dimensional problems due to slow convergence rate, and problems with a single mode as the samplers are prone to be trapped at a local mode. Annealed Importance Sampler (AIS, Neal (2001)) and Sequential Monte Carlo (SMC, (Del Moral et al., 2006)) were developed for sampling challenging multi-modal distributions and are considered the state-of-the-art MC methods for such problems. Variational Inference (VI, Wainwright & Jordan (2008)) is an alternative approach, which turns the sampling problem into an optimization one by fitting an easy-to-sample parametric distribution to $\nu$, often by minimizing the reverse (or exclusive) Kullback–Leibler (KL) divergence. The choice of the reverse KL is commonly thought to be due to computational convenience but can be validated by an information theoretical argument of optimal information processing of Bayes’ theorem (Zellner, 1988). For VI, it is popular to adopt mean-field approximation (Wainwright & Jordan, 2008) or a Normalizing Flow (VI-NF, Rezende & Mohamed (2015)) as the modeling parametric distribution. More recently, a plethora of neural network (NN) based algorithms, for example, neural network gradient HMC (Li et al., 2019), Annealed Flow Transport Monte Carlo (AFTMC, Arbel et al. (2021a)), Path-Integral Sampler (PIS, Zhang & Chen (2022)), adaptive MC with NF Gabrié et al. (2022), Continual Repeated AFTMC (CR-AFTMC, Matthews et al. (2022)), and Denoising Diffusion Sampler (DDS, (Vargas et al., 2023a)). For estimating the marginalized likelihood, simple MCMC methods require augmented algorithms (e.g., Green (1995)) or additional processing of the derived samples (e.g., Robert & Wraith (2009); Pajor (2017)). SIS and SMC can estimate the marginalized likelihood and are considered the gold-standard MC estimates. Among the NN-based algorithms, AFTMC, CR-AFTMC, PIS, and DDS are capable of estimating the marginalized likelihood. Llorente et al. (2023) provide a more comprehensive review of methods estimating the marginalized likelihood.
@@ 2.2 Learning velocity field for sampling
We remark that our methodology significantly differs from the Flow Matching approach (Lipman et al., 2023; Tong et al., 2024). Flow Matching trains a continuous-time normalizing flow by matching it to a known velocity field. This technique is specifically designed for training Generative Models using samples of a data distribution. In contrast, our task involves learning the normalizing flow using a prescribed path of annealed distributions, $\tilde{\rho}_{\ast}(x,t)$. In this scenario, we do not have access to the samples of the target distribution nor a target velocity field for matching: the problem would have been solved if we already had the samples of the target distribution, and we could have simply driven the samples by a known target velocity field if it is already known. Instead, our knowledge is limited to the fact that the velocity field at each of the samples must satisfy Eq. (5). Our goal is to solve for the unknown $v(x,t)$ in Eq. (5), rather than attempting to match a target flow.
@@ 3 Numerical Experiments
After training, we performed 30 independent samplings (each with $S=2000$ samples, different from those used in training) using the learned velocity fields. Following the recent series of papers (Arbel et al., 2021a; Zhang & Chen, 2022; Vargas et al., 2023a), we focus on assessing sample qualities by the estimation of log normalization constant $\log\hat{\mathcal{Z}}$. For low dimensional ($D\leq 10$) problems, we also compare the sliced Wasserstein distance $W_{p}$ (Bonneel et al., ) to ground-truth samples with $p=2$ as an additional metric. For all the methods with importance sampling, we used the weighted samples for computing these statistical quantities.
@@ 3.1 Testing problems
Mode-separated Gaussian mixture (type-1): An illustrative model of $D=2$ Gaussian mixture distribution with separated modes. We consider a similar challenging distribution as in PIS (Zhang & Chen, 2022): a mixture of nine Gaussian distributions centered at the grid $\{-1,0,1\}^{2}$, and each Gaussian has variance 0.012. We present results with equally-weighted ($1/9$) modes in the main text, and unequally-weighted modes in Appendix D.2.
@@ 3.2 Results
Tables 1 and 2 show the performance of different methods on type-1 and type-2 problems respectively. We visualize the samples and weight distributions of type-1 problems in Fig. 2. Comprehensive results of different experimental settings can be found in Appendix D. Our numerical results showed that LFIS is capable of generating competitive samples to existing methods. For two type-1 applications, we observed reasonable estimates of $\log\hat{\mathcal{Z}}$. Ground-truth (GT) samples can be generated for these two test problems, and the samples generated by LFIS have the lowest sample-to-GT-sample Wasserstein-2 distance. We remark that LFIS performs even better than the conventionally regarded gold-standard SMC with the same number of scales ($T=256$). For Effective Sample Size (ESS) (Liu, 2008), LFIS also shows competitive performance among all methods studied. Note that both SMC and AFT perform resampling when ESS is lower than a certain threshold (0.98 for SMC and 0.3 for AFT (Arbel et al., 2021a)), while LFIS, DDS, PIS do not resample. Even without resampling, LFIS can still achieve high ESS comparable to SMC with a high resampling threshold. The samples generated by LFIS also provide more uniform coverage over the different modes of the Gaussian mixture and the tails of funnel distribution compared to all other methods. In addition, the weight distributions of LFIS in these two problems are significantly narrower than other methods, explaining the higher ESS. For type-2 problems, LFIS is the only method delivering estimates of $\log\hat{\mathcal{Z}}$ that are consistent with the gold-standard SMC (with $T=1024$) on all problems.
@@ 4 Discussion
VI-NF (Rezende & Mohamed, 2015) uses a normalizing flow as the modeling distribution, which is parametrized by variational inference. LFIS shares many similar features of VI-NF when a Neural ODE (Chen et al., 2018) is used as the flow. In this case, both our proposition and VI-NF draw samples from simple initial distributions ($\mu$ in our case; the base distribution in VI-NF) and use the NN-modeled flow to transport the samples to a future time. Both methods leverage the computable density function along the sample trajectory, a unique feature of flow-based models (see Eq. (8)). The key difference between our proposition and VI-NF is the learning target. VI-NF is an end-to-end approach, that the loss function only depends on matching the final distribution at the end of the flow to the target distribution. In contrast, LFIS aims to capture $\tilde{\rho}_{\ast}(t)$, $t\in\left[0,1\right]$. We hypothesized that providing the whole evolution $\tilde{\rho}_{\ast}(t)$, instead of only the final $\tilde{\nu}=\tilde{\rho}_{\ast}(1)$, could improve the learning [^3] As the same philosophy applies to AIS and SMC (analogous to LFIS) to simple MCMC-type inference (analogous to VI-NF).. Moreover, VI-NF aims to directly optimize the reverse Kullback–Leibler (KL) divergence. In contrast, our approach first establishes the PDE (5) for the optimal flow which perfectly matches the evolution of the target distribution. Then, we parametrize the neural flow model by minimizing its discrepancy to the evolutionary equation, an approach similar to the Physics Informed Neural Networks (Raissi et al., 2019; Karniadakis et al., 2021). Because VI-NF has been shown to perform worse than AFTMC (Arbel et al., 2021a), we did not include VI-NF in our quantitative comparison in Sec. 3.
@@
We conclude the manuscript by listing the potential limitations of LFIS. (1) LFIS cannot handle non-differentiable density function $\tilde{\nu}(x)$. For example, bounded uniform priors. Theoretically, these singularities can be handled by inserting sources at the discontinuities, a direction that merits future developments. A practical solution to address the challenge is to consider the non-differentiable density function as the limit of differentiable $\tilde{\rho}_{\ast}(x,t)$. For example, using sigmoid functions $\tilde{\nu}(x,t)=\sigma(tx/(1-t))$ which converge to step functions $\tilde{\nu}(x)={\Theta}(x)$. (2) When the flow satisfying Eq. (5) is too complex, LFIS’ approach may require a more expressive NN than end-to-end DDS and PIS require. This can result in a more resource-intensive training process, but we speculate that it might offer a tradeoff in terms of improved accuracy. (3) LFIS is more memory-demanding as it requires $\nabla_{x}S_{\ast}$ in the equation, and cross terms like $\partial_{\theta}\partial_{x}v_{\theta}(x,t)$ in the optimization step. These terms are currently evaluated by memory-demanding autograd. (4) As an integrator and with the currently adopted explicit scheme, LFIS does not perform well when $T$ is small due to the error induced by a finite time-step (see Fig. 7). This may result in a higher training cost. Optimizing LFIS using higher-order integration schemes and an interpolation of the neural flow along the time domain merits future research.
Code availability
The code for LFIS and the results of numerical experiments have been deposited at https://github.com/lanl/LFIS.
Acknowledgments
The authors acknowledge continual support from Laboratory Directed Research and Development (LDRD). YT was supported by LDRD project “Accelerated Dynamics Across Computational and Physical Scales” (220063DR), NP was supported by LDRD project “Learning Uncertainties In Coupled-Physics Models via Operator Theory” (20230254ER), and YTL was supported by LDRD project “Diffusion Modeling with Physical Constraints for Scientific Data” (20240074ER). YTL sincerely thanks Prof. A. Doucet for several insightful email exchanges that partially inspired this work.
Impact Statement
This manuscript presents work whose goal is to advance the field of statistical sampling and machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix B Importance Sampling for Imperfect Neural Networks
We put a stringent condition that $\rho_{\theta}$ has to dominate $\rho_{\ast}$ for any arbitrary function $F$. The condition is often relaxed to “$\rho_{\theta}(x)>0$ where $F(x)\rho_{\ast}(x)>0$” in the importance sampling literature (e.g., Tokdar & Kass (2010)), which requires knowledge of the function $F$.
@@ D.1 Additional details of the test distributions
For the sampling problem in the latent space of VAE, we took the decoder $p_{\theta}(x|z)$ of the pre-trained VAE in Arbel et al. (2021a) and used it to construct the likelihood function with binary cross-entropy loss. In Fig. 3, we demonstrate the test image and the reconstructed image using the pre-trained decoder $p_{\theta}(x|z)$ from samples generated by LFIS.
@@ D.3 Neural network architecture
It is difficult to make a direct comparison of NN structures and numbers of parameters between the discrete-time and continuous-time models. For LFIS, the total number of parameters will increase with the total number of steps. For DDS and PIS, the number of parameters will increase with the number of channels used for time-encoding. Overall, when LFIS uses the same number of time steps as the time-encoding in DDS and PIS, the total number of NN parameters are very similar, thus causing no additional memory footprint.
@@ D.4 Additional details of the training procedure
The general training procedure is summarized in Algorithm 1. For each discretized time step $k$, at each training epoch, a new minibatch (B) of samples is generated by calling the GenerateSamples function. This step is crucial to provide good coverage of the sampling space. However, for low-dimensional problems, repeatedly calling the GenerateSamples function for each minibatch could slow down the training process, because the initial samples need to be passed through all the previously trained flow $v_{\theta_{\ast}^{(l)}}$. To accelerate the training of low-dimensional sampling problems ($D<10$), we skip the generation of new samples at each epoch. Instead, we use a subset of the $N$ samples used for estimating $\langle\partial_{t}\log\tilde{\rho_{\ast}}\rangle_{\ast}$ for training. When $N$ is large enough to provide good coverage of the sampling space, this simplified training procedure could provide a noticeable acceleration of learning the Liouville flow.
@@ D.9 Choice of total time steps and tempered scales
$T$ MG ($D=2$) Funnel ($D=10$) VAE ($D=30$)
32 $-0.064\pm 0.065$ $-0.43\pm 0.050$ $-220.70\pm 2.23$
$\widehat{\log\mathcal{Z}}$ 64 $-0.019\pm 0.060$ $-0.27\pm 0.066$ $-110.58\pm 0.39$
128 $-0.007\pm 0.066$ $-0.23\pm 0.057$ $-110.07\pm 0.31$
256 $0.004\pm 0.059$ $-0.11\pm 0.075$ $-109.95\pm 0.25$
32 $0.006\pm 0.005$ $-0.29\pm 0.023$ $-215.64\pm 0.87$
$\log\hat{\mathcal{Z}}$ 64 $0.002\pm 0.003$ $-0.16\pm 0.028$ $-109.98\pm 0.01$
128 $-0.0001\pm 0.004$ $-0.20\pm 0.011$ $-109.95\pm 0.01$
256 $-0.0002\pm 0.004$ $-0.07\pm 0.003$ $-109.99\pm 0.08$
$T$ Ionosphere ($D=35$) Sonar ($D=61$) LGCP ($D=1600$)
32 $-112.28\pm 0.68$ $-133.36\pm 2.08$ $470.13\pm 5.57$
$\widehat{\log\mathcal{Z}}$ 64 $-112.04\pm 0.38$ $-108.94\pm 0.30$ $489.31\pm 3.20$
128 $-111.67\pm 0.26$ $-108.60\pm 0.31$ $505.26\pm 5.56$
256 $-111.63\pm 0.33$ $-108.48\pm 0.30$ $505.43\pm 3.00$
32 $-111.42\pm 0.01$ $-130.45\pm 0.233$ $463.40\pm 2.61$
$\log\hat{\mathcal{Z}}$ 64 $-111.58\pm 0.006$ $-108.36\pm 0.011$ $484.54\pm 2.01$
128 $-111.59\pm 0.004$ $-108.37\pm 0.008$ $504.32\pm 1.56$
256 $-111.60\pm 0.006$ $-108.38\pm 0.009$ $505.53\pm 0.95$